# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v2

@@ Acknowledgment
We are deeply grateful to Ion Victor Gosea at the Max Planck Institute for Dynamics of Complex Technical Systems in Magdeburg, Germany, for his patient responses to our numerous questions about the Loewner framework and for his valuable feedback. We are also thankful to Prof. Patrick Kürschner at Leipzig University of Applied Sciences for his patient responses to our numerous questions about ADI method. This work is supported by the National Natural Science Foundation of China under Grants No. 62350410484 and 62273059, and in part by the High-end Foreign Expert Program No. G2023027005L granted by the State Administration of Foreign Experts Affairs (SAFEA).

### v3

@@ 6 Conclusion
Acknowledgements We are deeply grateful to Ion Victor Gosea at the Max Planck Institute for Dynamics of Complex Technical Systems in Magdeburg, Germany, for his patient responses to our numerous questions about the Loewner framework and for his valuable feedback. We are also thankful to Prof. Patrick Kürschner at Leipzig University of Applied Sciences for his patient responses to our numerous questions about the ADI method.
Statements and Declarations
Funding
This work is supported by the National Natural Science Foundation of China under Grants No. 62350410484 and 62273059.
Competing Interests
The authors declare no competing interests.
Consent for publication
All authors have read and approved the final manuscript.
Data availability
The MATLAB code and data to reproduce the results of this paper are publicly available at [62].
Authors’ Contributions
Umair Zulfiqar developed the main results of the paper and wrote the first manuscript. Qiu-Yan Song, Zhi-Hua Xiao, and Victor Sreeram contributed equally by providing critical improvements to the mathematical results and the manuscript’s presentation. Their contributions significantly enhanced the final draft. They also validated the MATLAB code to ensure the reproducibility of the numerical results.